OpenStudy (igreen):

The mitochondrial DNA sequence that is shared by two species has a steady mutation rate. Scientists determine the sequences of these two species to be as follows: Species A: CAGGCCATTATG Species B: CCAGCCTATAGG This DNA sequence has a known mutation rate, which the scientists used to calculate that the species diverged from a common ancestor 80 million years ago. Using this information, how much more time do you predict will pass before these species differ by a total of eight base pairs? 160 million years 32 million years 48 million years 240 million years

OpenStudy (igreen):

I keep coming up with 128. Because 80 million years / 5 differences = 16. 16 * 8 = 128

OpenStudy (igreen):

@ganeshie8

ganeshie8 (ganeshie8):

@Jhannybean

OpenStudy (igreen):

Lol

OpenStudy (jhannybean):

Oh this is like permutation kind of problem. No?

OpenStudy (igreen):

I guess so.

OpenStudy (jhannybean):

base pairs are like G-C, A-T, etc. And clearly they're not transcribed into U or anything.

OpenStudy (jhannybean):

So I would look for the base pairs first, and perhaps because they repeat at every interval or so.... (only a guess), that to predict 80 years you need to find where the pattern starts repeating once again. Again, this is just an opinion, lol.

OpenStudy (igreen):

Um, okay...?

OpenStudy (igreen):

So how do you find the Base Pairs?

OpenStudy (igreen):

So not all the letters in the sequence are base pairs?

OpenStudy (igreen):

Any ideas?

OpenStudy (jhannybean):

Oh I kind of get it.

OpenStudy (jhannybean):

So, I was on the right track sort of. First you need to count the number of base changes that happen in the sequence.

OpenStudy (jhannybean):

Let me point them out, one sec.

OpenStudy (igreen):

Yes, that's 5.

OpenStudy (igreen):

\(Species~A \! \!: C\color{red}A\color{red}GGCCA\color{red}TTA\color{red}TG\) \(Species~B \!\!: C\color{red}C\color{red}AGCCT\color{red}ATA\color{red}GG\)

OpenStudy (igreen):

Oh..that's four..lol.

OpenStudy (igreen):

No, wait that is 5..xd

OpenStudy (jhannybean):

OK yeah I was trying to type that out in \(LaTeX\) but I couldnt.

OpenStudy (igreen):

\(Species~A \! \!: C\color{red}A\color{red}GGCC\color{red}A\color{red}TTA\color{red}TG\) \(Species~B \!\!: C\color{red}C\color{red}AGCC\color{red}T\color{red}ATA\color{red}GG\)

OpenStudy (jhannybean):

Im so bad at it <.< but anyways..

OpenStudy (igreen):

Lol

OpenStudy (igreen):

All 5 differences aren't base pairs, are they?

OpenStudy (jhannybean):

So you already have 5 changes, and you can predict the initial (or current) change which is \(\frac{5}{80} = \frac{ 1\ \ mut}{16 \ \ years}\)

OpenStudy (igreen):

Yes, I got that far.

OpenStudy (jhannybean):

1 mutation happens every 16 years, and you need 3 more.

OpenStudy (jhannybean):

What do you do?

OpenStudy (igreen):

\(\dfrac{5}{80} = \dfrac{ 1\ \ mut}{16 \ \ years}\) I did 5 differences / 80 million years = 1 difference 16 million years So 8 different base pairs * 16 million years = 128 million years. But 128 million years isn't an option.

OpenStudy (jhannybean):

no but see we're already have done 5 different base pairs 8-5 = 3 pairs left. 3 * 16 million years = 48 million years.

OpenStudy (jhannybean):

That's how much time is left for all 8 to occur within the time frame of 80 years, (our approx)

OpenStudy (igreen):

Oh..

OpenStudy (igreen):

Thanks a bunch.

OpenStudy (jhannybean):

No problem :)

OpenStudy (praetorian.10):

oh yeah this totally interests me, remind me to never click any link you send me on skype again xD

OpenStudy (igreen):

..............